Apple Bite's is a small distributor of Ernest Inc. in a highly competitive healthcare industry. The market determining price of their product is $10. If apple Bite's TC function is TC= 100+ 2Q + 0.10 Q².

a) what is the firm's profit maximizing output?

aii) what is the firm's profit at the firm's profit maximizing output level?

iii) is apple in the long run or short run equilibrium? Explain

b) at $10, what is apple Bite's break even output level?

c) what is apple Bite's long run, break even price and output level?

d) what is apple Bite's shutdown price and output level? Does these price and output combination constitute the long run or short run competitive equilibrium. Explain



1
Expert's answer
2021-04-07T10:21:30-0400

(a) The firm maximizes its profit when "P=MC":


"10=2+0.2Q,""Q=40."

(aii) We can find the firm's profit as follows:


"\pi=TR-TC=PQ-100-2Q-0.1Q^2,""\pi=\$10\cdot40-100-2\cdot40-0.1\cdot(40)^2=\$60."